# Beyond Matryoshka: Revisiting Sparse Coding for Adaptive Representation

### v2

@@ 6 Conclusion
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.1 Evaluation Metric
We adopt 1-NN as our evaluation metric, implemented using FAISS (Johnson et al., 2019) with exact L2 search, following the setup in (Kusupati et al., 2022). This approach provides an efficient and cost-effective way to evaluate the utility of learned representations for downstream tasks, as 1-NN accuracy requires no additional training. In detail, we use the training set with 1.3M samples as the database and the validation set with 50K samples as the query set.
@@ B.3 Implementation Detail
We select the ResNet50 model [^7] https://huggingface.co/timm/resnet50d.ra4_e3600_r224_in1k as our backbone from (Wightman, 2019). For image preprocessing, we adopt the same procedure as described in (Kusupati et al., 2022; Leclerc et al., 2023). Consistent with (Gao et al., 2024), we utilize a tied encoder-decoder structure to build the CSR framework. The implementation of CSR is based on the codebase [^8] https://github.com/openai/sparse_autoencoder provided by OpenAI. All experiments are conducted on a server equipped with 4 RTX4090 GPUs. The selection of hyperparameters are:
@@
Reset50 2048 512 4e-5 10 1024 512 1/32 1.0 8,16,32…2048 Adam 1e-4 6.25 * 1e-10
Table 3: Implementation details on Image experiment.
@@
1-NN classification results is shown in B.4. The Int8 quantization operation is performed using a sentence transformer on pre-trained ResNet50 embeddings.
Table 4: 1-NN results of different methods on ImageNet1k classification.
Active Dim MRL MRL-E SVD Rand. LP SAE CSR Quant Int8 Golden Performance
8 62.19 57.45 24.61 6.06 69.69 73.84 - -
16 67.91 67.05 51.02 16.96 71.51 74.39 - -
32 69.46 68.60 64.98 36.99 72.90 74.53 - -
64 70.17 69.61 71.24 51.37 74.12 74.62 - -
128 70.52 70.12 73.63 62.53 74.28 74.78 - -
256 70.62 70.36 74.77 67.32 74.36 74.86 - -
512 70.82 70.74 75.08 72.17 74.38 74.88 - -
1024 70.89 71.07 75.14 74.49 74.40 74.91 - -
2048 70.97 71.21 75.19 75.19 74.41 74.96 73.48 75.19
@@ C.3 Baselines
We choose several models that provide MRL embeddings on MTEB benchmark (Muennighoff et al., 2022). These models are Stella-en-1.5B-v5 (Zhang et al., 2025), Jina-V3 (Sturua et al., 2024), Nomic-Embed-V1.5 (Nussbaum et al., 2024), Gecko-Text-Embedding-004-256 (Lee et al., 2024b), OpenAI-Text-Embedding-3-L-256 (OpenAI, 2024), Arctic-Embed-L-V2.0 (Yu et al., 2024), Potion-Base-2M (min, 2024).
@@ C.4 Implementation Detail
We select NV-EmbedV2 (Lee et al., 2024a) as our pre-trained model. We utilize a tied encoder-decoder structure to build the CSR framework. For text classification and clustering tasks, we use data from the same class as positive samples while the other as negative samples to calculate Equation 5. The hyperparameters are set as follows:
@@
NV-EmbeV2 4096 16384 4e-5 10 128 1024 0.1 1.0 32,64,256 Adam 1e-4 6.25 * 1e-10
Table 5: Implementation details on Text experiment
@@ D.2 Experiment Setup
We selected ViT-B-16, trained on the DFN2B dataset [^9] https://huggingface.co/apple/DFN2B-CLIP-ViT-B-16, as our pre-trained model. Following (Kusupati et al., 2022), we implemented MRL in the pre-trained ViT model and fine-tuned it for 50 epochs on the MSCOCO (Lin et al., 2014) and Flickr30K (Young et al., 2014) datasets, respectively. For a fair comparison, we also fine-tuned the backbone on both datasets for 50 epochs using the same hyperparameters, which were then used for the backbone of CSR . The hyperparameters used for fine-tuning are as follows:
@@
Table 6: Hyperparameters for fine-tuning ViT-B/16 backbone.
@@ D.3 Implementation Detail
We select the ViT-B-16101010https://huggingface.co/apple/DFN2B-CLIP-ViT-B-16as our backbone from (Wightman, 2019). Consistent with (Gao et al., 2024), we utilize a tied encoder-decoder structure to build the CSR framework. The encoder and decoder structure share between image space and text space. The implementation of CSR is based on the codebase [^11] https://github.com/openai/sparse_autoencoder and OpenCLIP (Cherti et al., 2023). The metric is evaluated through CLIP-benchmark following standard procedure. All experiments are conducted on a server equipped with 4 RTX4090 GPUs. We present detailed training parameters for the multimodal experiment in Table 7.
@@
Table 7: Implementation details on MultiModal experiment.
D.4 Disscussion On Dead Latents
Addressing the mitigation of dead latents in the alignment space remains an open challenge, leaving room for future work and study. For a detailed analysis, please refer to Section D.4. Table 2 presents the performance comparison between CSR and MRL, revealing that the gap between the two methods diminishes as sparsity constraints become more stringent. Further analysis indicates that CSR continues to face the “dead latents” issue despite incorporating advanced mechanisms. As shown in Figure 8, CSR exhibits a significant performance drop, corresponding to a sharp rise in dead latent dimensions. We attribute this to a technical challenge, as CSR has demonstrated robust performance in both image and text domains under similar sparsity constraints. This suggests that representations in alignment spaces may require more specialized design, presenting an opportunity for future research and improvement.
@@ E.2 Effect on Hidden Representation Dimension $\mathbb{R}^{h}$
Implementation details are shown in Table 9. The pre-trained ViT-L/16141414https://huggingface.co/timm/vit_large_patch16_224.augreg_in21k_ft_in1k and ResNet50 models [^15] https://huggingface.co/timm/resnet50.a1_in1k can be found at timm (Wightman, 2019). To ensure generalizability, we train the model using three different random seeds and report the mean performance in the main paper.

### v4

@@ 6 Conclusion & Discussion
Our method, CSR, is orthogonal to existing acceleration techniques such as pruning (He et al., 2017), quantization (Jacob et al., 2018), and distillation (Hinton et al., 2015), which primarily target embedding generation. In contrast, CSR optimizes the post-processing stage, enabling complementary speedups with minimal performance trade-off. A current limitation of CSR, shared by other sparsity-based approaches, is the emergence of dead neurons under high sparsity, especially in multimodal settings. While techniques like contrastive loss partially mitigate this (see Figure 6), fully resolving the issue remains an open challenge and direction for future work.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant U21B2006; in part by the Fundamental Research Funds for the Central Universities QTZX24003 and QTZX23018; in part by the 111 Project under Grant B18039; and in part by Shaanxi Youth Innovation Team Project.
@@ Acknowledgements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.1 Evaluation Metric
We adopt 1-NN as our evaluation metric, implemented using FAISS (Johnson et al., 2019) with exact L2 search, following the setup in (Kusupati et al., 2022). This approach provides an efficient and cost-effective way to evaluate the utility of learned representations for downstream tasks, as 1-NN accuracy requires no additional training. In detail, we use the training set with 1.3M samples as the database and the validation set with 50K samples as the query set. We also report linear probing and few-shot results using Top-1 accuracy. For a holistic evaluation, different methods, Figure 1 (c) presents the average 1-NN performance (active dimensions $<64$).
@@ B.3 Implementation Detail
For a fair comparison, we selected the pre-trained ResNet50 weights, noted as FF2048 in the MRL  (Kusupati et al., 2022). Additionaly, we select the ResNet50 model [^5] https://huggingface.co/timm/resnet50d.ra4_e3600_r224_in1k as our SOTA backbone from  Wightman (2019). For image preprocessing, we adopt the same procedure as described in  Kusupati et al. (2022); Leclerc et al. (2023). Consistent with  Gao et al. (2024), we utilize a tied encoder-decoder structure to build the CSR framework. The implementation of CSR is based on the codebase [^6] https://github.com/openai/sparse_autoencoder provided by OpenAI. All experiments are conducted on a server equipped with 4 RTX4090 GPUs. The selection of hyperparameters are:
Table 3: Implementation details on Image experiment.
@@
ResNet50 2048 8192 4e-5 10 4096 512 1/32 0.1 8,16,32…2048 Adam 1e-4 6.25 * 1e-10
@@
1-NN classification and Top-1 linear probing results are shown in Table 4 and Table 5.
Table 4: 1-NN accuracy of different methods on ImageNet1k classification.
Active Dim Full Rep. MRL MRL-E SVD Rand. FS SAE CSR SOTA Full Rep. CSR (w/ SOTA RN50)
8 - 62.19 57.45 19.14 2.36 67.14 67.78 - 73.84
16 - 67.91 67.05 46.02 12.06 68.14 69.17 - 74.39
32 - 69.46 68.60 60.78 32.91 68.91 70.15 - 74.53
64 - 70.17 69.61 67.04 49.91 69.69 70.94 - 74.62
128 - 70.52 70.12 69.63 60.91 69.74 70.99 - 74.65
256 - 70.62 70.36 70.67 65.75 70.35 71.31 - 74.73
512 - 70.82 70.74 71.06 68.77 71.21 71.29 - 74.88
1024 - 70.89 71.07 71.22 70.41 71.20 71.30 - 74.90
2048 71.19 70.97 71.21 71.21 71.19 71.24 71.20 75.19 74.91
Table 5: Top-1 classification accuracy results of different methods on ImageNet1k classification.
Active Dim Full Rep. MRL MRL-E MRL-LP SVD Rand. LP SAE CSR SOTA Full Rep. CSR (w/ SOTA RN50)
8 - 66.63 56.66 5.15 2.34 4.56 73.46 73.62 - 79.17
16 - 73.53 71.94 13.79 7.17 11.29 74.60 74.75 - 79.72
32 - 75.03 74.48 32.52 20.46 27.21 75.28 75.44 - 79.96
64 - 75.82 75.35 52.66 48.10 49.47 75.81 75.88 - 80.16
128 - 76.30 75.80 64.60 67.24 65.70 75.91 76.24 - 80.24
256 - 76.47 76.22 69.29 74.59 72.43 76.27 76.25 - 80.31
512 - 76.65 76.36 70.51 76.78 74.94 76.43 76.34 - 80.33
1024 - 76.76 76.48 70.19 76.87 76.10 76.59 76.54 - 80.32
2048 76.87 76.80 76.51 69.72 - 76.87 76.66 76.52 80.59 80.35
@@ C.3 Baselines
We choose several models that provide MRL embeddings on MTEB benchmark (Muennighoff et al., 2022). These models are Stella-en-1.5B-v5 (Zhang et al., 2025), Jina-V3 (Sturua et al., 2024), Nomic-Embed-V1.5 (Nussbaum et al., 2024), Gecko-Text-Embedding-004-256 (Lee et al., 2024b), OpenAI-Text-Embedding-3-L-256 (OpenAI, 2024), Arctic-Embed-L-V2.0 (Yu et al., 2024) and Potion-Base-2M (min, 2024).
@@ C.4 Implementation Detail
We select NV-Embed-V2 (Lee et al., 2024a) as our pre-trained model. We utilize a tied encoder-decoder structure to build the CSR framework. For text classification and clustering tasks, we use data from the same class as positive samples while the other as negative samples to calculate Equation 6. The hyperparameters are set as follows:
Table 6: Implementation details on Text experiment.
@@
NV-Embed-V2 4096 16384 4e-5 10 128 1024 0.1 1.0 32,64,256 Adam 1e-4 6.25 * 1e-10
@@ D.2 Experiment Setup
We selected ViT-B-16, trained on the DFN2B dataset [^7] https://huggingface.co/apple/DFN2B-CLIP-ViT-B-16, as our pre-trained model. For the in-distribution cross-modal retrieval experiment, we implemented MRL in the pre-trained ViT model following  Kusupati et al. (2022), and fine-tuned it for 50 epochs on the MSCOCO (Lin et al., 2014) and Flickr30K (Young et al., 2014) datasets, respectively. For a fair comparison, we also fine-tuned the backbone on both datasets for 50 epochs using the same hyperparameters, which were then used for the backbone of CSR . The hyperparameters used for fine-tuning are as follows:
Table 7: Hyperparameters for fine-tuning ViT-B/16 backbone.
@@
For zero-shot cross-modal retrieval, we employed the same MRL fine-tuning procedure as in our in-distribution experiment, maintaining identical hyperparameters while training for 3 epochs with 2208 batch size on CC3M (Changpinyo et al., 2021).
@@ D.3 Implementation Detail
We select the ViT-B-16888https://huggingface.co/apple/DFN2B-CLIP-ViT-B-16as our backbone from  Wightman (2019). Consistent with  Gao et al. (2024), we utilize a tied encoder-decoder structure to build the CSR framework. The encoder and decoder structure share between image space and text space. The implementation of CSR is based on the codebase [^9] https://github.com/openai/sparse_autoencoder and OpenCLIP (Cherti et al., 2023). The metric is evaluated through CLIP-benchmark following standard procedure. All experiments are conducted on a server equipped with 4 RTX4090 GPUs. We present detailed training parameters for the multimodal experiment in Table 8.
Table 8: Implementation details on MultiModal experiment.
@@
CC3M 512 4096 4e-4 1 1024 1024 1/32 1.0 64,128,256 Adam 0.0 6.25 * 1e-10
D.4 Discussion On Dead Latents
Addressing the mitigation of dead latents in the alignment space remains an open challenge, leaving room for future work and study. Table 2 presents the performance comparison between CSR and MRL, revealing that the gap between the two methods diminishes as sparsity constraints become more stringent. Further analysis indicates that CSR continues to face the “dead latents” issue despite incorporating advanced mechanisms. As shown in Figure 8, CSR exhibits a significant performance drop, corresponding to a sharp rise in dead latent dimensions. We attribute this to a technical challenge, as CSR has demonstrated robust performance in both image and text domains under similar sparsity constraints. This suggests that representations in alignment spaces may require more specialized design, presenting an opportunity for future research and improvement.
@@ E.2 Effect on Hidden Representation Dimension $\mathbb{R}^{h}$
Implementation details are shown in Table 10. The pre-trained ViT-L/16121212https://huggingface.co/timm/vit_large_patch16_224.augreg_in21k_ft_in1k and ResNet50 models [^13] https://huggingface.co/timm/resnet50.a1_in1k can be found at timm (Wightman, 2019). To ensure generalizability, we train the model using three different random seeds and report the mean performance in the main paper.
@@ E.3 Retrieval Time Evaluation
E.4 Understanding Retrieval Time Difference between Dense and Sparse Embeddings
Although CSR and MRL have similar theoretical complexity $O(k)$, their actual runtimes are affected by backend implementations. For instance, cuBLAS (used for dense ops) is highly optimized but has high launch overhead, while cuSPARSE (used for CSR) is lighter but less optimized for small $k$. Here, we can share a preliminary insight into why sparse embeddings can be faster than dense embeddings and why it can get faster with larger hidden dimension $h$.
Sparse matrix multiplication benefits from zero-skipping: only overlapping non-zero entries are used. For each query, computing the $i$-th output only involves comparing indices of non-zero entries—an integer operation much cheaper than floating-point multiplication. As $h$ increases and $k$ stays small, overlap likelihood drops, reducing the number of multiplications required. In Table We empirically verify this by counting the number of multiplications under various $h$:
Table 11: Comparison on the number of multiplication operation between MRL (dense) and CSR (embeddings) on the default setup.
Active Dim MRL CSR ($h=8192$) CSR ($h=16384$) CSR ($h=32768$)
2 $1.3\times 10^{9}$ $3.2\times 10^{5}$ $1.7\times 10^{5}$ $8.4\times 10^{4}$
4 $2.6\times 10^{9}$ $1.3\times 10^{6}$ $6.7\times 10^{5}$ $3.4\times 10^{5}$
The number of operations in CSR is several orders of magnitude smaller than in MRL, and it decreases with larger $h$. This counterintuitive yet practical effect highlights the appeal of using sparse high-dimensional embeddings: they allow richer representations while improving runtime.